Ten android tablets will be delivered to a certain school for testing but three of them are defective. The principal will get two of these tablets for testing. Let X represent the defective tablet, Y represent the non-defective tablet and L represent the random variable representing the defective tablet will occur. Complete the table below and find the values of random variable L.


1
Expert's answer
2021-03-29T12:59:49-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 \text{Possible } & \text{Values of the } \\
 \text{Outcomes} & \text{Random Variable} \ L\\
\hline
LL & 2 \\
LY & 1 \\
YL & 1\\
YY & 0

\end{array}"

"P(L=0)=\dfrac{\dbinom{3}{0}\dbinom{10-3}{2-0}}{\dbinom{10}{2}}=\dfrac{21}{45}=\dfrac{7}{15}"

"P(L=1)=\dfrac{\dbinom{3}{1}\dbinom{10-3}{2-1}}{\dbinom{10}{2}}=\dfrac{21}{45}=\dfrac{7}{15}"

"P(L=2)=\dfrac{\dbinom{3}{2}\dbinom{10-3}{2-2}}{\dbinom{10}{2}}=\dfrac{3}{45}=\dfrac{1}{15}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 l & 0 & 1 & 2 \\ 
 \hdashline \\

 P(L=l) & \dfrac{7}{15} & \dfrac{7}{15} & \dfrac{1}{15}
\end{array}"